# MLPs Learn In-Context

### v1

@@ 2.2 ICL classification
Following Reddy [9], we present in-context classification as follows. The input consists of a sequence of context exemplars $(\boldsymbol{x}_{1},\boldsymbol{y}_{1}),(\boldsymbol{x}_{2},\boldsymbol{y}_{2}),\ldots,(\boldsymbol{x}_{L},\boldsymbol{y}_{L})$ followed by a query point $\boldsymbol{x}_{q}$, where $\boldsymbol{x}_{i},\boldsymbol{y}_{i}\in\mathbb{R}^{n}$. The $\boldsymbol{x}$ points are sampled from a Gaussian mixture model $\mathcal{M}_{k}$ consisting of $k$ components. Each mixture component (i.e. cluster) is labeled by one among $C$ labels, where $k\geq C$, so multiple clusters may map to the same label. Labels are represented in the context by vectors $\boldsymbol{\alpha}_{1},\boldsymbol{\alpha}_{2},\ldots\boldsymbol{\alpha}_{C}\in\mathbb{R}^{n}$. If $\boldsymbol{x}_{i}$ belongs to cluster $j$, then $\boldsymbol{y}_{i}=\boldsymbol{\alpha}_{j}$. The model must predict the correct label for $\boldsymbol{x}_{q}$, and outputs $C$ logits corresponding to the $C$ labels (not a vector of values $\boldsymbol{\alpha}$, which are used only to represent labels in the context). Figure 1g illustrates this task, with additional details in Appendix C.7.
@@ Results.
We first compare how different architectures perform at ICL across different compute in Figure 1h. The task is parameterized by burstiness $B=4$ and $k=2048$ with $L=8$ points in the context, a setting in which all models develop an in-context solution (see Figure 6d for details). As before, with sufficient compute Transformers do not outperform vanilla MLPs or Mixers. Indeed, at small compute, vanilla MLPs attain a markedly lower loss. Note: in this setting, there are $L/B=2$ labels present in each context, out of $C=32$ total possible labels. As a baseline, we plot in gray the loss obtained by placing equal probability on the 2 labels present in-context. MLPs in particular appear to plateau at this baseline before approaching zero loss with higher compute.
We also measure how well each model handles increasingly long context lengths in Figure 1i. In a surprising reversal, vanilla MLPs attain a relatively flat loss across context lengths, as do Transformers. Mixers’ loss increases somewhat for longer contexts. Overall, MLPs continue to perform at a comparable level with Transformers on in-context classification.
Finally, we observe a transition from IWL to ICL across the three architectures as data diversity increases. As in Reddy [9], we measure IWL by constructing test examples where the query point does not appear in the context. The only way the model correctly classifies these points is if it memorizes the mapping from cluster to label. To measure ICL, we consider two different strategies: 1) sample points from an entirely different mixture $\mathcal{M}_{k}^{\prime}$, producing novel clusters, or 2) swap cluster/label mappings so that clusters are labeled differently than they were during training. Test examples from either strategy can only be solved if the model identifies cluster labels in-context, since the underlying cluster label assignment is different from training.  [^3] In practice, accuracy on these two ICL measures is virtually identical across all models and settings. We plot accuracy across all three types of test examples in Figure 1j for increasing $k$, and observe a transition from IWL to ICL across all three model architectures. The transition happens for somewhat lower data diversity in Transformers and Mixers, and somewhat higher in vanilla MLPs. Additional plots and parameterizations are explored in Appendix D.
@@ 3 Experiment: Relational tasks
In addition to the classic regression and classification paradigms, we also consider tasks from cognitive science used to study the capacity for relational reasoning in humans, animals, and neural networks [17, 18, 19, 30]. These tasks are a subset of in-context classification, and rely on understanding distance relations between context exemplars. Many of the original tasks use visual stimuli. To remove confounds associated with learning sufficiently rich visual representations and clarify the role of relational reasoning, we provide real-valued inputs that function as idealized stimulus embeddings.
In a step up from the general ICL tasks explored in Section 2, we find that MLPs perform better with less compute than Transformers, and generalize more effectively on out-of-distribution test sets. To compare against a benchmark for gold-standard performance using hand-crafted features, we consider a relationally bottlenecked MLP (RB MLP), an architecture that has been demonstrated to solve a variety of challenging relational tasks with competitive generalization and efficiency characteristics [31, 17]. Relational bottlenecks are architectural components that prevent absolute information about the inputs from propagating downstream; rather, the RB computes a set of (hand-crafted) relations — often simple dot products — between inputs, and allows only these relations to flow downstream, forcing the model to operate at a higher level of abstraction. Our RB MLP operates by first computing dot product relations between inputs, then propagating optionally through several MLP layers before a final readout (see Appendix C.5 for details). We find that while relational bottlenecks are helpful when the model’s relations align well with the task structure, they may fail on tasks with deviating structure. All in all, these relational tasks demonstrate that MLPs can non-trivially outperform Transformers in a controlled setting, suggesting potential practical advantage for their usage.
The question of whether neural network models can reason relationally at all has been an enduring topic of heated debate (see Alhama and Zuidema [32] for a recent review). Our results fall decisively in favor of yes, they can, and contrast a recent attempt at formally proving that MLPs cannot reason relationally [33]. In Appendix A, we discuss our relationship with the relational reasoning literature further, comment on the proof by Boix-Adsera et al. [33], and demonstrate empirically that, under the right conditions, a vanilla MLP solves a task posited by their arguments to be impossible.
@@ 3.1 Match to sample
The match-to-sample (MTS) task paradigm originates in Skinnerian behavioral experiments [18]. A test subject is first presented with a sample stimulus (e.g. an image). The subject is then shown a set of many stimuli, and must select the one that matches the original sample.
@@
In addition to the vanilla MLP and Transformer models, we also consider a relationally bottlenecked MLP architecture (RB MLP) [31]. The RB MLP uses dot-product relations between the query point and each of the five context points $\boldsymbol{r}=\left(\boldsymbol{x}_{q}\cdot\boldsymbol{x}_{1},\boldsymbol{x}_{q}\cdot\boldsymbol{x}_{2},\ldots,\boldsymbol{x}_{q}\cdot\boldsymbol{x}_{L}\right)$. The relations $\boldsymbol{r}$ are passed directly to a softmax readout, producing class predictions $\boldsymbol{y}_{\text{RB}}=\mathsf{smax}\left(\boldsymbol{W}_{\text{readout}}\boldsymbol{r}\right)$. Note, a choice of weights $\boldsymbol{W}_{\text{readout}}=\boldsymbol{I}$ solves the task perfectly, though it remains to be seen whether the RB model discovers this solution. Further details on the RB MLP model are in Appendix C.5.
@@ Results.
Figure 2b plots the loss obtained by each of the three models on the MTS task as a function of compute. Surprisingly, the vanilla MLP outperforms the Transformer by a notable margin. With relations that are well-aligned to the task, the RB MLP model achieves the best compute efficiency.
We also consider how well each model performs in different kinds of out-of-distribution test examples. Results are plotted in Figure 2c. We first try perturbing the radius of the circle after training with unit radius. As we vary the radius during testing, both MLP models continue to perform well, but the Transformer quickly degenerates. We also try changing the order of context points. If the points are ordered, they are presented in clockwise order along the circle. If the points are scrambled, they are presented in random order. Curiously, if the models are trained first on ordered points, then no model generalizes well when subsequently tested with scrambled points. This is true even for the relationally bottlenecked model, indicating that it fails to learn the full generalizing solution $W_{\text{readout}}=I$ when presented with an ordered context.
@@ 3.2 Sphere oddball
The oddball tasks described in the next two sections are based on work from Sablé-Meyer et al. [19], who used it to measure geometric relational reasoning in humans and baboons. In an oddball task, the test subject is presented with six stimuli, originally abstract geometric images. One image differs from the others. The subject must select this “oddball” to succeed. Like the MTS task, the oddball tasks are a subset of ICL classification where all-but-one point belong to the same class.
As before, our versions of the oddball task simplify full visual stimuli into point embeddings, representing a pre-learned image representation. The model is presented with $L$ context points $\boldsymbol{x}_{1},\boldsymbol{x}_{2},\ldots,\boldsymbol{x}_{L}\in\mathbb{R}^{2}$. (There are no query points.) In the sphere oddball task, the context points are sampled as $\boldsymbol{x}\sim\mathcal{N}(\boldsymbol{\mu},\boldsymbol{I})$, for some nonzero center $\boldsymbol{\mu}$. One point in the context is randomly selected and perturbed in a random direction by a distance $d$. The model must return the index of this oddball point. In the results that follow, we use $L=6$ points in the context. Figure 2d presents an illustration of the task, with additional details in Appendix C.9.
@@ 4 Discussion
Overall, we observed that MLPs can learn in-context and moreover, perform at a level comparable to Transformers on in-context tasks. We also examined relational reasoning tasks, a subset of ICL classification, which have historically been considered too challenging for simple connectionist architectures like MLPs [32, 34]. Surprisingly, MLPs learn these relational tasks well — and exhibit both better compute efficiency and generalization performance than Transformers! This observation challenges earlier claims [33, 31], but is consistent with the emerging realization that, given sufficient data diversity and compute, an MLP can indeed learn to reason relationally [30]. We discuss our relationship with prior work in relational reasoning further in Appendix A.
Broadly, our work is consistent with the “bitter lesson" of AI research [13]: in the face of increasing compute and data resources, general methods with weaker inductive bias will outperform methods endowed with stronger inductive bias. This heuristic speaks to the intuition that a strong inductive bias may be beneficial for particular tasks, but may misalign the model in different or more general settings. We see an extreme example of this in studying relationally bottlenecked MLPs, where hand-crafted relations strongly benefit the model in specific cases where they align with the task. However, departing even slightly from the ideal task structure prevents the shallow RB MLP from learning the task at all, while a vanilla MLP continues to exhibit strong performance. In the absence of hand-designed relations, Transformers are more general learners than RB MLPs, but less than vanilla MLPs. As a result, for certain well-suited tasks (like ICL regression), Transformers perform measurably more efficiently for a fixed compute budget. But for other tasks (relational reasoning, simple regression and classification in Appendix B), MLPs have the upper hand.
These results expand the range of possible tasks commonly thought solvable by MLP models. ICL is clearly not the exclusive domain of Transformers, and we encourage greater engagement with ICL beyond attention-based architectures. The surprising success of MLPs for relational reasoning also encourages a change in perspective about how simple connectionist architectures may be capable of solving relational tasks, and under what conditions they fail. The impressive performance of MLPs hints at potential real-world benefits, and we watch the future outcomes of all-MLP approaches with interest.
@@
We consider only controlled, simple tasks designed to probe specific characteristics of each model. We never compare architectures on complex datasets like those found in language or vision, though there are other studies that do, and find that MLPs continue to perform competitively [12, 20, 21]. The advantage of our approach is that we avoid confounds irrelevant to ICL introduced by complex data, and clarify the specific competencies of each model to learn in-context across representative settings. Nonetheless, an important direction for future work is to study how MLPs perform on more complex tasks, and characterize failure modes if they occur.
@@
Overall, we hope our results encourage further work studying ICL beyond attention-based architectures, and the properties of simple connectionist architectures like MLPs that enable them to solve relational tasks. Important questions remain in quantifying how much data diversity is generally required to transition to ICL, how this threshold depends on architecture, varying sensitivity to context length across architectures, precisely characterizing differences in inductive bias for ICL, and more. As compute and data resources continue to grow year-over-year, we watch with interest whether less inductive bias will indeed be better.
Acknowledgements
We thank Alex Atanasov, Hamza Chaudhry, Alex Meterez, Mo Osman, Sab Sainathan, Jacob Zavatone-Veth, and members of the Pehlevan Group for many helpful comments and discussions on our manuscript. WLT was supported by a Kempner Graduate Fellowship. CP and WLT were supported by NSF Award DMS-2134157 and NSF CAREER Award IIS-2239780. CP is further supported by a Sloan Research Fellowship. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence. The computations in this paper were run on the FASRC cluster supported by the FAS Division of Science Research Computing Group at Harvard University.
@@ Appendix A MLPs reason relationally
Relational reasoning is closely related to in-context learning, as solving ICL tasks requires reasoning about the relationships between inputs while ignoring their absolute characteristics. Indeed, the relational tasks we explore in the main text are specific subsets of ICL classification. At the same time, MLPs are commonly presumed to fail at relational reasoning [34, 35]. While our primary focus remains on comparing in-context learning between Transformers and MLPs, we offer this digression to contextualize our results within the broader relational reasoning literature.
The question of whether connectionist models are able to reason relationally at all has been an enduring topic of passionate debate (see Alhama and Zuidema [32] for a recent review). Our empirics support the notion that classic architectures like vanilla MLPs can indeed reason relationally, consistent with recent findings in Geiger et al. [30]. However, many researchers presuppose that classic architectures cannot solve relational tasks, resulting in a zoo of alternative architectures aimed at endowing neural networks with relational capacities [31, 36, 30, 32].
One especially strong claim that MLPs cannot reason relationally was advanced by Boix-Adsera et al. [33], who formally prove that MLPs will never generalize to unseen symbols on relational tasks. Their proof however relies on an unconventional input scheme that hinders learning. Below, we discuss their analysis on MLPs and offer our own remarks on the learnability of relational tasks. We also demonstrate empirically that, under a more conventional setting, MLPs do generalize on a relational task claimed by Boix-Adsera et al. [33] to be impossible.
@@
A concrete example of this setup is the same-different task often used in probing relational reasoning [30]. Here, we consider two templates $z_{1}=\alpha\alpha$ and $z_{2}=\alpha\beta$. The labeling function is $f^{*}(z_{1})=1$ and $f^{*}(z_{2})=0$. The string $\boldsymbol{x}=AA$ matches template $z_{1}$, and the string $\boldsymbol{x}^{\prime}=AB$ matches template $z_{2}$. We will abuse notation slightly and write $f^{*}(\boldsymbol{x})=f^{*}(\boldsymbol{z}_{1})=1$. Crucially, matching a template depends only on relations between symbols and not on the symbols themselves. For example,
@@
Their proof relies on the permutation invariance property of MLPs and SGD [37]. Summarizing briefly, they argue that if $x_{1},x_{2}\in\mathcal{X}_{uns}$, we can swap their one-hot encodings without any impact on the MLP. More generally, we can construct a permutation matrix $\Pi\in\mathbb{R}^{k|\mathcal{X}|\times k|\mathcal{X}|}$ such that for all strings of symbols $\boldsymbol{x}^{(1)},\boldsymbol{x}^{(2)}\in\mathcal{X}_{uns}^{k}$ and $\boldsymbol{x}^{\prime}\in\mathcal{X}_{seen}^{k}$, it remains true that $\Pi\,\mathsf{flat}(\boldsymbol{E}(\boldsymbol{x}^{(1)}))=\mathsf{flat}(\boldsymbol{E}(\boldsymbol{x}^{(2)}))$ and $\Pi\,\mathsf{flat}(\boldsymbol{E}(\boldsymbol{x}^{\prime}))=\mathsf{flat}(\boldsymbol{E}(\boldsymbol{x}^{\prime}))$. That is, we permute only the indices of unseen symbols, but leave the indices of seen symbols untouched. Then given the permutation symmetry of MLPs and SGD [37], because we preserve the indices of the seen symbols, we must have that

### v3

@@ 2.2 ICL classification
Following Reddy (2024), we present in-context classification as follows. The input consists of a sequence of context exemplars $(\boldsymbol{x}_{1},\boldsymbol{y}_{1}),(\boldsymbol{x}_{2},\boldsymbol{y}_{2}),\ldots,(\boldsymbol{x}_{L},\boldsymbol{y}_{L})$ followed by a query point $\boldsymbol{x}_{q}$, where $\boldsymbol{x}_{i},\boldsymbol{y}_{i}\in\mathbb{R}^{n}$. The $\boldsymbol{x}$ points are sampled from a Gaussian mixture model $\mathcal{M}_{k}$ consisting of $k$ components. Each mixture component (i.e. cluster) is labeled by one among $C$ labels, where $k\geq C$, so multiple clusters may map to the same label. Labels are represented in the context by vectors $\boldsymbol{\alpha}_{1},\boldsymbol{\alpha}_{2},\ldots\boldsymbol{\alpha}_{C}\in\mathbb{R}^{n}$. If $\boldsymbol{x}_{i}$ belongs to cluster $j$, then $\boldsymbol{y}_{i}=\boldsymbol{\alpha}_{j}$. The model must predict the correct label for $\boldsymbol{x}_{q}$, and outputs $C$ logits corresponding to the $C$ labels (not a vector of values $\boldsymbol{\alpha}$, which are used only to represent labels in the context). Figure 1g illustrates this task, with additional details in Appendix C.7.
@@ Results.
We first compare how different architectures perform at ICL across different compute in Figure 1h. The task is parameterized by burstiness $B=4$ and $k=2048$ with $L=8$ points in the context, a setting in which all models develop an in-context solution (see Figure 7d for details). As before, with sufficient compute Transformers do not outperform vanilla MLPs or Mixers. Indeed, at small compute, vanilla MLPs attain a somewhat lower loss. Note: in this setting, there are $L/B=2$ labels present in each context, out of $C=32$ total possible labels. As a baseline, we plot in gray the loss obtained by placing equal probability on the 2 labels present in-context. MLPs in particular appear to plateau at this baseline before approaching zero loss with higher compute.
We also measure how well each model handles increasingly long context lengths in Figure 1i. In a surprising reversal, vanilla MLPs attain a relatively flat loss across context lengths, as do Transformers. Mixers’ loss increases somewhat for longer contexts, though this blip vanishes at higher dimensions (Figure 7b). Overall, MLPs continue to perform at a comparable level with Transformers on in-context classification.
Finally, we observe a transition from IWL to ICL across the three architectures as data diversity increases. As in Reddy (2024), we measure IWL by constructing test examples where the query point does not appear in the context. The only way the model correctly classifies these points is if it memorizes the mapping from cluster to label. To measure ICL, we consider two different strategies: 1) sample points from an entirely different mixture $\mathcal{M}_{k}^{\prime}$, producing novel clusters, or 2) swap cluster/label mappings so that clusters are labeled differently than they were during training. Test examples from either strategy can only be solved if the model identifies cluster labels in-context, since the underlying cluster label assignment is different from training. [^3] In practice, accuracy on these two ICL measures is virtually identical across all models and settings. We plot accuracy across all three types of test examples in Figure 1j for increasing $k$, and observe a transition from IWL to ICL across all three model architectures. The transition happens for somewhat lower data diversity in Transformers and Mixers, and somewhat higher in vanilla MLPs. Additional plots and parameterizations are explored in Appendix D.
@@ 3 Experiment: Relational tasks
We next consider classical tasks from psychology used to study relational reasoning in humans, animals, and neural networks (Campbell et al., 2023; Skinner, 1950; Sablé-Meyer et al., 2021; Geiger et al., 2023). These tasks are functionally a subset of in-context classification, and rely on understanding similarity relationships between context exemplars.
In a surprising advance from the tasks explored in Section 2, we find that MLPs perform better with less compute than Transformers, and generalize more effectively on out-of-distribution test sets. As a benchmark for gold-standard performance using hand-crafted features, we implement a relationally bottlenecked MLP (RB MLP), an architecture demonstrated to solve many challenging relational tasks with competitive generalization and efficiency characteristics (Webb et al., 2023; Campbell et al., 2023). Relational bottlenecks are architectural components that prevent absolute information about the inputs from propagating downstream; rather, the RB computes a set of (hand-crafted) relations between inputs (often simple dot products), and allows only these relations to flow downstream, forcing the model to operate on abstractions. Our RB MLP operates by first computing dot product relations between inputs, then propagating optionally through several MLP layers before a final readout (see Appendix C.5 for details). We find that while relational bottlenecks are helpful when the model’s relations align well with the task structure, they may fail on tasks with deviating structure. All in all, these relational tasks demonstrate that MLPs can quite surprisingly outperform Transformers on certain in-context tasks.
The question of whether neural network models can reason relationally at all has been an enduring topic of heated debate (see Alhama and Zuidema (2019) for a recent review). Our results fall decisively in favor of the affirmative, and contrast a recent attempt at formally proving that MLPs cannot reason relationally (Boix-Adsera et al., 2023). In Appendix A, we discuss our relationship with the relational reasoning literature, comment on the proof by Boix-Adsera et al. (2023), and demonstrate empirically that a vanilla MLP solves a task posited by their arguments to be impossible.
@@ 3.1 Match to sample
The match-to-sample (MTS) task paradigm originates in Skinnerian behavioral experiments (Skinner, 1950). A test subject is first presented with a sample stimulus (e.g. an image). The subject is then shown a set of many stimuli, and must select the one that matches the original sample.
@@
In addition to the vanilla MLP and Transformer models, we also consider a relationally bottlenecked MLP architecture (RB MLP) (Webb et al., 2023). The RB MLP uses dot-product relations $\boldsymbol{r}$ between the query point and each of the five context points $\boldsymbol{r}=\left(\boldsymbol{x}_{q}\cdot\boldsymbol{x}_{1},\boldsymbol{x}_{q}\cdot\boldsymbol{x}_{2},\ldots,\boldsymbol{x}_{q}\cdot\boldsymbol{x}_{L}\right)$. The relations $\boldsymbol{r}$ are passed directly to a softmax readout, producing class predictions $\boldsymbol{y}_{\text{RB}}=\mathsf{smax}\left(\boldsymbol{W}_{\text{readout}}\,\boldsymbol{r}\right)$. Note, a choice of weights $\boldsymbol{W}_{\text{readout}}=\boldsymbol{I}$ solves the task perfectly, though it remains to be seen whether the RB model discovers this solution. Further details on the RB MLP model are in Appendix C.5.
@@ Results.
Figure 2b plots the loss obtained by each of the three models on the MTS task as a function of compute. The vanilla MLP outperforms the Transformer by a surprising margin. With relations that are well-aligned to the task, the RB MLP model achieves the best compute efficiency.
We also consider how well each model performs in different kinds of out-of-distribution test examples. Results are plotted in Figure 2c. We first try perturbing the radius of the circle after training with unit radius. As we vary the radius during testing, both MLP models continue to perform well, but the Transformer quickly degenerates. We also try changing the order of context points. If the points are ordered, they are presented in clockwise order along the circle. If the points are scrambled, they are presented in random order. Curiously, if the models are trained first on ordered points, then no model generalizes well when subsequently tested with scrambled points (not even the relationally bottlenecked model).
@@ 3.2 Sphere oddball
The oddball tasks described in the next two sections are based on work from Sablé-Meyer et al. (2021), who used it to measure geometric relational reasoning in humans and baboons. In an oddball task, the test subject is presented with six stimuli, originally geometric images. One image differs from the others. The subject must select this “oddball” to succeed. Like the MTS task, the oddball tasks are a subset of ICL classification where all-but-one point belong to the same class.
As before, our version of the oddball task simplify full visual stimuli into idealized stimulus representations. The model is presented with $L$ context points $\boldsymbol{x}_{1},\boldsymbol{x}_{2},\ldots,\boldsymbol{x}_{L}\in\mathbb{R}^{2}$. (There are no query points.) In the sphere oddball task, the context points are sampled as $\boldsymbol{x}\sim\mathcal{N}(\boldsymbol{\mu},\boldsymbol{I})$, for some nonzero center $\boldsymbol{\mu}$. One point in the context is randomly selected and perturbed in a random direction by a distance $d$. The model must return the index of this oddball point. In the results that follow, we use $L=6$ points in the context. Figure 2d presents an illustration of the task, with additional details in Appendix C.9.
@@ 4 Discussion
We observed that MLPs can learn in-context and moreover, perform at a level comparable to Transformers on in-context tasks. We also examined relational reasoning tasks, closely related to ICL classification, which have historically been considered beyond the ability of simple neural architectures like MLPs (Alhama and Zuidema, 2019; Marcus, 1998; Boix-Adsera et al., 2023). Surprisingly, MLPs learn these relational tasks well — and exhibit both better compute efficiency and generalization performance than Transformers. This observation diverges from earlier claims (Boix-Adsera et al., 2023; Webb et al., 2023), but is consistent with the emerging realization that, given sufficient data diversity and compute, an MLP can indeed learn to reason relationally (Geiger et al., 2023). We discuss our relationship with prior work in relational reasoning further in Appendix A.
Broadly, our work is consistent with the “bitter lesson” of AI research (Sutton, 2019): in the face of increasing compute and data resources, general methods with weaker inductive bias will outperform specialist methods endowed with stronger inductive bias. This heuristic speaks to the intuition that a strong inductive bias may be beneficial for particular tasks, but may misalign the model in different or more general settings. We see an extreme example of this in studying relationally bottlenecked MLPs, where hand-crafted relations strongly benefit the model in specific cases where they align with the task. However, departing even slightly from the ideal task structure prevents the shallow RB MLP from learning the task at all, while a vanilla MLP continues to exhibit strong performance. In the absence of hand-designed relations, Transformers are more general learners than RB MLPs, but less than vanilla MLPs. As a result, for certain well-suited tasks (like ICL regression), Transformers perform more efficiently for a fixed compute budget. But for other tasks (relational reasoning, simple regression and classification in Appendix B), MLPs have the upper hand.
These results expand the range of possible tasks commonly thought solvable by MLP models. ICL may not be the exclusive domain of Transformers, and we encourage greater engagement with ICL beyond attention-based architectures. The surprising success of MLPs for relational reasoning also encourages a change in perspective about how simple architectures may be capable of solving relational tasks, and under what conditions they fail. The impressive performance of MLPs hints at potential real-world benefits, and we watch the future outcomes of all-MLP approaches with interest.
@@
We consider only controlled, synthetic tasks designed to probe for specific characteristics. We never compare architectures on complex datasets like those found in language or vision, though there are other studies that do, and find that MLPs continue to perform competitively (Tolstikhin et al., 2021; Liu et al., 2021; Fusco et al., 2022). The advantage of our approach is that we avoid confounds irrelevant to ICL introduced by complex data, and clarify the specific competencies of each model to learn in-context across representative settings. It remains overall unclear how MLP-based architectures perform against Transformers at scale in complex, real-world tasks, and where a possible performance gap may originate. Future work should explore MLPs in more naturalistic settings to better understand the potential comparative advantage of attention-based schemes.
@@
Overall, we hope our results encourage further work studying ICL beyond attention-based architectures, and the properties of simple architectures like MLPs that enable them to solve relational tasks. Important questions remain in quantifying how much data diversity is generally required to transition to ICL, how this threshold depends on architecture, varying sensitivity to context length across architectures, precisely characterizing differences in inductive bias for ICL, and more.
Ethics statement.
Since this study focuses on synthetic tasks, it is limited in direct negative societal impacts beyond that of general theoretical machine learning research. We do not conduct experiments in human or animal subjects.
Reproducibility statement.
Full descriptions of all tasks are provided in Sections 2 and 3 of the main text. Additional details regarding specific task and model configurations, along with code availability, compute requirements, and software, are comprehensively enumerated in Appendix C.
Acknowledgements.
We thank Alex Atanasov, Hamza Chaudhry, Alex Meterez, Mo Osman, Sab Sainathan, Jacob Zavatone-Veth, members of the Pehlevan Group, and the anonymous ICLR reviewers for many helpful comments and discussions on our manuscript. WLT is supported by a Kempner Graduate Fellowship. CP is supported by NSF grant DMS-2134157, NSF CAREER Award IIS-2239780, DARPA grant DIAL-FP-038, a Sloan Research Fellowship, and The William F. Milton Fund from Harvard University. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence. The computations in this paper were run on the FASRC cluster supported by the FAS Division of Science Research Computing Group at Harvard University.
@@ Appendix A MLPs reason relationally
Relational reasoning is closely related to in-context learning, as solving ICL tasks requires reasoning about the relationships between inputs while ignoring their absolute characteristics. Indeed, the relational tasks we explore in the main text are functional subsets of ICL classification. At the same time, MLPs are commonly presumed to fail at relational reasoning [Marcus, 1998, Boix-Adsera et al., 2023] or exhibit severe weaknesses [Fodor and Pylyshyn, 1988]. While our primary focus remains on comparing in-context learning between Transformers and MLPs, we offer this digression to contextualize our results within the broader relational reasoning literature.
The question of whether connectionist models are able to reason relationally at all has been an enduring topic of passionate debate (see Alhama and Zuidema [2019] for a recent review). Our empirics support the notion that classical architectures like vanilla MLPs can indeed reason relationally, consistent with recent findings in Geiger et al. [2023]. However, many researchers presuppose that classical architectures cannot solve relational tasks, resulting in a zoo of alternatives aimed at endowing neural networks with relational capacities [Webb et al., 2023, Battaglia et al., 2018, Geiger et al., 2023, Alhama and Zuidema, 2019].
One especially strong claim that MLPs cannot reason relationally was advanced by Boix-Adsera et al. [2023], who formally prove that MLPs will never generalize to unseen symbols on relational tasks. Their proof however relies on a pathological input scheme that hinders learning. Below, we discuss their analysis on MLPs and offer our own remarks on the learnability of relational tasks. We also demonstrate empirically that, under conventional settings, MLPs do generalize on a relational task claimed by Boix-Adsera et al. [2023] to be impossible.
@@
A concrete example of this setup is the same-different task often used in probing relational reasoning [Geiger et al., 2023]. Here, we consider two templates $z_{1}=\alpha\alpha$ and $z_{2}=\alpha\beta$. The labeling function is $f^{*}(z_{1})=1$ and $f^{*}(z_{2})=0$. The string $\boldsymbol{x}=AA$ matches template $z_{1}$, and the string $\boldsymbol{x}^{\prime}=AB$ matches template $z_{2}$. We will abuse notation slightly and write $f^{*}(\boldsymbol{x})=f^{*}(\boldsymbol{z}_{1})=1$. Crucially, matching a template depends only on relations between symbols and not on the symbols themselves. For example,
@@
Their proof relies on the permutation invariance property of MLPs and SGD [Ng, 2004]. Summarizing briefly, they argue that if $x_{1},x_{2}\in\mathcal{X}_{uns}$, we can swap their one-hot encodings without any impact on the MLP. More generally, we can construct a permutation matrix $\Pi\in\mathbb{R}^{k|\mathcal{X}|\times k|\mathcal{X}|}$ such that for all strings of symbols $\boldsymbol{x}^{(1)},\boldsymbol{x}^{(2)}\in\mathcal{X}_{uns}^{k}$ and $\boldsymbol{x}^{\prime}\in\mathcal{X}_{seen}^{k}$, it remains true that $\Pi\,\mathsf{flat}(\boldsymbol{E}(\boldsymbol{x}^{(1)}))=\mathsf{flat}(\boldsymbol{E}(\boldsymbol{x}^{(2)}))$ and $\Pi\,\mathsf{flat}(\boldsymbol{E}(\boldsymbol{x}^{\prime}))=\mathsf{flat}(\boldsymbol{E}(\boldsymbol{x}^{\prime}))$. That is, we permute only the indices of unseen symbols, but leave the indices of seen symbols untouched. Then given the permutation symmetry of MLPs and SGD [Ng, 2004], because we preserve the indices of the seen symbols, we must have that